Evaluate |-30|x|1/5-1/3|. Thank you for your help!
Answer by tutorcecilia(2144)

|-30|x|1/5-1/3|
I would break this problem down into two parts.
First part: the absolute value of -30 is 30
Second part: the absolute value of 1/5-1/3 = |-2/15|=2/15
So (30)(2/15) = 4